A Gunn oscillator creates a resonant frequency of 24.15 GHz. What is the time of flight in nanoseconds within this oscillator? (Type in a four-decimal number)
1
Expert's answer
2020-07-16T07:03:24-0400

Resonant frequency is given by Fmax = 1/2td

Therefore,

given,

Fmax = 24.15*109 = 1/2td

td = 2.07039*10-11 sec

= 0.0207 nano seconds